OpenStudy (anonymous):

A camping stove, burning butane, was used to heat 500g of water from 20C until it was boiling. Heating this amount of water from 20C to boiling with an electrical heater requires 168Kj of energy. a) If the pot is made of aluminium and weighed 100g, how much heat energy was required to heat the pot (specific heat capacity of Al= 875jkgK) b) what is the total energy required to heat the pot and water? c) when the water started to boil, the stove weighed 14.5g less than it had initially. How many moles of butane were used to heat the pot and water?

sam (.sam.):

Q=mcT of water is 168KJ

sam (.sam.):

I don't think Q=mcT applies to aluminium

OpenStudy (anonymous):

Energy required to heat the pot: q=mcDeltaT Don't forget to convert 875 j/kg*K to 0.875 j/g*K q = 0.875 * 100 g Al * 80 K =7000 J required b.) q= 4.18 J/gK *500 g H2O * 80 K = 167,200 J c. C4H10 is butane. Molar mass is 58.12 g/mol. If the stove weighed 14.5 g less, then that means 14.5 grams of butane was burned off. 14.5 g C4H10 * ( 1 mol/58.12 g C4H10) = 0.25 mol butane

OpenStudy (anonymous):

Oops, for part B, I forgot to add the heat to warm up the aluminum pot. 167,200 J is the heat needed to heat the water. 7000 K is the heat needed to heat the pot. 167,200 J + 7000 J = 174,200 J.

OpenStudy (anonymous):

Thank you so so much!! I have no words to thank you! You are incredible!!!! :)

OpenStudy (anonymous):

No problem!